OpenStudy (anonymous):

use lagrange multipliers to find the extreme values of f(x,y)=x^2-y^2 subject to the constraint x^2+y^2=1. I) the extreme values of f(x,y), subject to said constraint, occur at those points (a,b) on R^2 such that <2a,-2b>=lambda<2a,2b> for some lambda on R and a^2+b^2=1. true or false?

OpenStudy (anonymous):

i answered true. im pretty sure it fits the first criteria. just not sure about the second. i assume the second one is by extension true

OpenStudy (dumbcow):

im not sure about the (a,b), sounds true if you work it out though the extreme value of f(x,y) =-1 for x=0, y=+-1

OpenStudy (dumbcow):

or y=0, x=+-1 works too

OpenStudy (anonymous):

does there need to be one value of lambda that satisfies the lagrange equation? I wrote out the lagrange equation and got: \[\Lambda(x,y,\lambda)=x^2-y^2+\lambda(x^2+y^2-1)\] taking partials: \[\Lambda_{x}=2x+2x \lambda=0, \Lambda_{y}=-2y+2y \lambda=0\] so for the first partial lambda must be negative, and for the second lambda must be positive. So there isn't a single multiplier so that the gradient of the lagrange equation is zero.

OpenStudy (anonymous):

i agree with it being true. the second part has three answer choices. a) the extrema occur at (+-1,0) and (0,+-1) b) the extreme values are -1 and 1 c) all of the above

OpenStudy (anonymous):

I geuss I need to study this more

OpenStudy (dumbcow):

thats what i got A) then

OpenStudy (anonymous):

what about the extreme values?

OpenStudy (dumbcow):

oops yeah so all of the above

OpenStudy (anonymous):

ok good. I feel better about this one too. I really appreciate the help.

OpenStudy (dumbcow):

rsvitale, i think its ok if multiplier can be either 1 or -1

OpenStudy (anonymous):

allright so we can have a seperate multiplier for x and y

OpenStudy (dumbcow):

as long as solution fits all three equations

OpenStudy (anonymous):

allright

OpenStudy (anonymous):

i appreciate the help from both of you.

OpenStudy (dumbcow):

your welcome, good luck with rest of your classes

OpenStudy (anonymous):

thanks